fara:

A sequence is defined by the recursive formula f (n + 1) = f(n) – 2. If f(1) = 18, what is f(5)?

sealake101:

ok so what have u learned abt this already?

fara:

thank you so much guys

justus:

Hello there. Here are the equations from 1 through 5. You can get the answer to f(5) by calculating it. f(1)=18 f(2)=f(1)-2=18-2= f(3)=f(2)-2=16-2= f(4)=f(3)-2=14-2= f(5)=f(4)-2=12-2=